# A Bias-Correction Decentralized Stochastic Gradient Algorithm with Momentum Acceleration

### v1

@@ 3.1 Drawbacks of momentum-based DSGD algorithm
Decentralized momentum SGD (DmSGD) (Yu et al., 2019; Gao and Huang, 2020) was proposed to accelerate the convergence of distributed learning. In DmSGD, each agent $i$ iteratively updates its momentum and communicates parameters, as follows:
@@
However Proposition 2 in Yuan et al. (2021) proved that DmSGD suffers from inconsistency bias for strongly convex loss under full-batch gradients (i.e. when $\sigma^{2}=0$),
@@
where $\mathbf{x}^{\star}$ is the optimum point, and $\zeta^{2}=(1/n)\sum_{i=1}^{n}\left\|\nabla f_{i}(\mathbf{x}^{*})-\nabla f(\mathbf{x}^{*})\right\|^{2}$ is defined as the measure of the data heterogeneity (Hsieh et al., 2020; Koloskova et al., 2020). Through the proof of theorem 3 in Yu et al. (2019), it can also be found that for the non-convex objective function $f_{i}(\mathbf{x})$, DmSGD will also converge to a neighborhood of local optimal, the radius of which is dominated by the variance of the stochastic gradient $\sigma^{2}$ and the data heterogeneity $\zeta^{2}$. When the data are more heterogeneous, DmSGD will oscillate over a larger area of the optimum, and the corresponding error will be larger. Therefore, how to use momentum to accelerate the decentralized algorithm while reducing heterogeneity is an urgent problem to be solved.
@@ 3.2 From Exact-Diffusion to Proposed Method
The challenge of heterogeneity is inherent in distributed SGD. Intuitively, we need to ensure that the parameter at the optimal value point still falls into the optimal position after one step of iteration. However, as long as the data is heterogeneous, that is, $\nabla f_{i}(\mathbf{X}^{\star})\neq 0$, where $\mathbf{X}^{\star}=\mathbf{1}{\mathbf{x}^{\star}}^{\top}$ and $\mathbf{x}^{\star}$ represents the optimal point, then the algorithm will shift the optimal value after one step gradient update $\mathbf{X}^{(t+1)}=\mathbf{X}^{\star}-\alpha\nabla\mathbf{f}(\mathbf{X}^{\star})\neq\mathbf{X}^{\star}$. Consequently, it is essential to employ specific methods to correct the biases introduced by the inconsistency in the gradient. Currently, a class of bias-correction distributed SGD algorithms is designed to address heterogeneity, which is summarized as the SUDA algorithm in Alghunaim and Yuan (2022). In this paper, we focus on a specific class of algorithms based on the ED/D2 framework, which can be summarized in the following three steps: first, perform local gradient descent represented by $\bm{\psi}_{i}^{(t+1)}=\mathbf{x}_{i}^{(t)}-\alpha\nabla{F}_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})$, where we use $\bm{\psi}_{i}^{(t)}$ to denote the temporary variable. Second, correct the bias caused by the data heterogeneity as expressed by $\bm{\phi}_{i}^{(t+1)}=\bm{\psi}_{i}^{(t+1)}+\mathbf{x}_{i}^{(t)}-\bm{\psi}_{i}^{(t)}$. Finally, combine the corrected parameter $\bm{\phi}_{i}^{(t+1)}$ to update the local parameter with $\mathbf{x}_{i}^{(t+1)}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\bm{\phi}_{j}^{(t+1)}$, where $\mathcal{N}_{i}=\{j:(i,j)\in\mathcal{E}\}$ represents the set of neighbors for agent $i$.
The main idea of ED/D2 is to use local historical information to correct the bias caused by the data heterogeneity in the local gradient update step. Thus ED/D2 will ultimately converge to a neighborhood of the optimum, being influenced solely by the randomness inherent in the stochastic gradient. Generally, the iteration of ED/D2 can be written as the following formula,
@@
To introduce the idea of momentum acceleration into ED/D2, we replace $\nabla F_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})$ with a weighted sum over the historical gradient, i.e. $\mathbf{m}_{i}^{(t)}$ of DmSGD in (3.2), and we get the Exact-Diffusion with Momentum (EDM) algorithm, which can be formally written in Algorithm 1. The iteration formula can be written as
Assume that $\mathbf{X}^{(t)}$ is obtained by Algorithm 1, and $\bar{\mathbf{x}}^{(t)}=\frac{1}{n}\mathbf{X}^{(t)}\mathbf{1}_{n}$. Then $\bar{\mathbf{x}}^{(t+1)}=\bar{\mathbf{x}}^{(t)}-\alpha\bar{\mathbf{m}}^{(t)}$. This implies that though the parameters are different on each agent, EDM has the same update pattern as SGDm in the average sense.
@@ 4.1 Auxiliary Variables
To measure the convergence rate, we are mainly concerned with $f(\bar{\mathbf{x}}^{(t)})$ and its gradient. We first define an auxiliary sequence for $\bar{\mathbf{x}}^{(t)}$, which is commonly used to analyze momentum-based SGD methods. Let $\mathbf{z}^{(0)}=\bar{\mathbf{x}}^{(0)}$ and
@@
The following two lemmas establish the connection between $\mathbf{z}^{(t)}$ and $\mathbf{X}^{(t)}$ in the non-convex and PL conditions, respectively.
@@ 4.2 Pseudo Deviation Variance Decomposition
According to Lemma 1 and 2, our next step is to conduct an in-depth investigation of $\mathbb{E}\|\mathbf{X}^{(t)}-\overline{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}$. To make the notation more concise, we introduce $\mathbf{P}_{\mathbf{I}}=\mathbf{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$, thus $\|\mathbf{X}^{(t)}-\overline{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}=\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\|_{\mathrm{F}}^{2}$.
The updating structure of momentum makes it difficult to exploit the independence of $\bm{\Xi}^{(t)}$ in the stochastic gradient. To deal with this difficulty, we introduce the following variables. Denote $\widetilde{\mathbf{M}}^{(-1)}=\mathbf{0}$ and $\widetilde{\mathbf{M}}^{(t)}=(1-\beta)\sum_{j=0}^{k}\beta^{t-j}\nabla\mathbf{f}(\mathbf{X}^{(j)})$, $\widetilde{\mathbf{X}}^{(-1)}=\widetilde{\mathbf{X}}^{(0)}=\mathbf{X}^{(0)}$ and
@@
Applying Cauchy’s inequality, we have
@@
Suppose that Assumptions 1 and 3 hold. And $\mathbf{X}^{(t)}$ is the sequence of parameters obtained from algorithm 1 and $\widetilde{\mathbf{X}}^{(t)}$ is defined above, then
@@
This lemma gives us an upper bound of $\mathbb{E}\|\mathbf{P}_{\mathbf{I}}(\mathbf{X}^{(t)}-\widetilde{\mathbf{X}}^{(t)})\|_{\mathrm{F}}^{2}$, which is not affected by $\beta$. Thus, we should only focus on the pseudo-deterministic part $\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}$. We say that it is pseudo-deterministic because it eliminates randomness in gradients, but still contains randomness in $\mathbf{X}^{(t)}$.
@@ 4.3 Transformations and Consensus Inequalities
In this part, we first show that $\widetilde{\mathbf{X}}^{(t)}$ has a recursive structure similar to SUDA in Alghunaim and Yuan (2022). Based on the ideas in Alghunaim and Yuan (2022), the updating formula can be rewritten as follows. Let $\widetilde{\mathbf{X}}^{(0)}=\mathbf{X}^{(0)}$, $\widetilde{\mathbf{M}}^{(0)}=(1-\beta)\nabla\mathbf{f}(\mathbf{X}^{(0)})$, $\widetilde{\mathbf{M}}^{(-1)}=\mathbf{0}$, $\widetilde{\mathbf{Y}}^{(0)}=\mathbf{0}$, for $t\geq 0$,
@@
The following lemma presents a transformation of $\widetilde{\mathbf{X}}^{(t)}$, which is used to find the upper bound of $\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{X}}^{(t)}\|^{2}$, and consequently for $\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\|^{2}$. This lemma establishes an upper bound on the data consensus deviation that will occur during iteration, which is central to our construction of the convergence properties of algorithms.
@@
Here, we do not specify the exact form of $\widetilde{\mathbf{N}}^{(t)}$. Because taking different values for $\widetilde{\mathbf{N}}^{(t)}$ in different proofs can simplify our proof, even though both are used to measure the difference between the current amount of updates and $\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t)})$. In the subsequent proofs, we will take $\widetilde{\mathbf{N}}^{(t)}=(1-\beta)\sum_{j=0}^{t}\beta^{t-j}\nabla\mathbf{f}(\overline{\mathbf{X}}^{(j)})+\beta^{t+1}\nabla\mathbf{f}(\overline{\mathbf{X}}^{(0)})$ for the non-convex case and $\widetilde{\mathbf{N}}^{(t)}=\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t)})$ for the PL condition.
@@ 4.4 Non-Convex
In practical optimization problems, we typically select $\beta\in[0.8,0.99)$, allowing us to treat $\beta$ as a fixed value. When $\beta=0$, the algorithm simplifies to the ED/D2 method. And Theorem 5 is consistent with the latest conclusions of ED/D2 in Alghunaim and Yuan (2022) with the corresponding bounds illustrated in Table 1. In contrast to their conclusions, we obtained a bound that is unaffected by the minimum eigenvalue of the communication matrix through a refined scaling of the initial values. Consequently, we arrive at a tighter upper bound when $\beta=0$. Consequently, the momentum-based algorithm exhibits a convergence rate that is at least comparable to the ED/D2 algorithm. This feature shows the strength of our algorithm.
@@
Unlike Quasi-Global and DecentLaM, which are also the momentum-based variants of DSGT, EDM can ultimately eliminate the effects of data heterogeneity. Compared with another kind of bias-correction algorithm (DSGT and DSGT-HB), EDM can eliminate data heterogeneity at a rate of $O(\alpha^{2}\zeta_{0}^{2}(1-\lambda)^{-2}T^{-1})$, while DSGT is $O(\alpha^{2}\zeta_{0}^{2}(1-\lambda)^{-3}T^{-1})$. This phenomenon will be shown in our subsequent real-data analysis.
@@ 4.5 PL-Condition
We compare EDM with other momentum-based distributed algorithms under the PL condition. Both of the convergence bounds of Quasi-Global and DecentLaM cannot eliminate the effect of heterogeneity. Currently, there is no work that proves the complete convergence properties of the DSGT algorithm with momentum under the Polyak-Lojasiewicz (PL) condition, or even under the strong convexity condition. Huang et al. (2024) demonstrates convergence properties when $\beta$ is sufficiently large. However, it does not extend to the original algorithm (that is, when $\beta=0$). In contrast, our findings impose no constraints on the choice of $\beta$ and cover the conclusion of ED/D2. Moreover, our approach can be generalized to prove convergence properties for momentum-based DSGT algorithms.
On the other hand, Huang et al. (2024) uses the technique of Loopless Chebyshev Acceleration (LCA) to reduce the communication gap in DSGT from $(1-\lambda)^{-1}$ to $(1-\lambda)^{-1/2}$. Thus, a better upper bound on convergence is obtained. However, this technique is not exclusive to DSGT, and other distributed algorithms can also reduce the communication gap by integrating LCA. This is not the focus of this article, so we will not compare it with such algorithms.

### v2

@@ 3.1 Drawbacks of momentum-based DSGD algorithm
Decentralized momentum SGD (DmSGD) (Yu et al., 2019; Gao and Huang, 2020) was proposed to accelerate the convergence of distributed learning. In DmSGD, each agent $i$ iteratively updates its momentum and communicates parameters as follows:
@@
However, Proposition 2 in Yuan et al. (2021) demonstrated that DmSGD suffers from an inconsistency bias for strongly convex loss under full-batch gradients (i.e. when $\sigma^{2}=0$),
@@
where $\mathbf{x}^{\star}$ denotes the optimum point, and $\zeta^{2}=(1/n)\sum_{i=1}^{n}\left\|\nabla f_{i}(\mathbf{x}^{*})-\nabla f(\mathbf{x}^{*})\right\|^{2}$ is defined as the measure of the data heterogeneity (Hsieh et al., 2020; Koloskova et al., 2020). Additionally, through the proof of theorem 3 in Yu et al. (2019), it can be shown that for the non-convex objective function $f_{i}(\mathbf{x})$, DmSGD will also converge to a neighborhood of local optimal, with the radius being influenced by the variance of the stochastic gradient $\sigma^{2}$ and the data heterogeneity $\zeta^{2}$. When the data exhibit greater heterogeneity, DmSGD tends to oscillate over a larger area surrounding the optimum, resulting in larger corresponding errors. Therefore, addressing the challenge of leveraging momentum to accelerate the decentralized algorithm while simultaneously reducing heterogeneity is an urgent problem that needs to be resolved.
@@ 3.2 From Exact-Diffusion to Proposed Method
The challenge of heterogeneity is inherent in distributed SGD. To ensure the stability of the convergence, we need to guarantee that the parameter at the optimal value point remains in the optimal position after one step of iteration. However, the heterogeneity of data commonly leads to the condition that $\nabla f_{i}(\mathbf{X}^{\star})\neq 0$ (where $\mathbf{X}^{\star}=\mathbf{1}{\mathbf{x}^{\star}}^{\top}$ and $\mathbf{x}^{\star}$ denotes the optimal point). Consequently, the algorithm will move away from the optimal value after one step gradient update, resulting in $\mathbf{X}^{(t+1)}=\mathbf{X}^{\star}-\alpha\nabla\mathbf{f}(\mathbf{X}^{\star})\neq\mathbf{X}^{\star}$. Consequently, it is essential to employ specific methods to correct the biases introduced by the inconsistency in the gradient. A class of bias-correction distributed SGD algorithms has been developed to address heterogeneity, summarized as the SUDA algorithm in Alghunaim and Yuan (2022). In this paper, we focus on a specific class of algorithms based on the ED/D2 framework, which can be summarized in the following three steps: first, perform local gradient descent represented by $\bm{\psi}_{i}^{(t+1)}=\mathbf{x}_{i}^{(t)}-\alpha\nabla{F}_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})$, where we use $\bm{\psi}_{i}^{(t)}$ to denote the temporary variable. Second, correct the bias caused by the data heterogeneity as expressed by $\bm{\phi}_{i}^{(t+1)}=\bm{\psi}_{i}^{(t+1)}+\mathbf{x}_{i}^{(t)}-\bm{\psi}_{i}^{(t)}$. Finally, combine the corrected parameter $\bm{\phi}_{i}^{(t+1)}$ to update the local parameter with $\mathbf{x}_{i}^{(t+1)}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\bm{\phi}_{j}^{(t+1)}$, where $\mathcal{N}_{i}=\{j:(i,j)\in\mathcal{E}\}$ represents the set of neighbors for agent $i$.
The main idea of ED/D2 is to utilize local historical information to correct the bias introduced by the data heterogeneity during the local gradient update step. Thus, ED/D2 will ultimately converge to a neighborhood of the optimum, influenced solely by the randomness inherent in the stochastic gradient. Generally, the iteration of ED/D2 can be expressed as follows:
@@
To incorporate the idea of momentum acceleration into ED/D2, we replace $\nabla F_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})$ with a weighted sum over the historical gradient, specifically $\mathbf{m}_{i}^{(t)}$ from DmSGD as defined in (3.2). This leads to the development of the Exact-Diffusion with Momentum (EDM) algorithm, which can be formally described in Algorithm 1. The iteration formula can be written as:
@@
Assume that $\mathbf{X}^{(t)}$ is obtained by Algorithm 1, and $\bar{\mathbf{x}}^{(t)}=\frac{1}{n}\mathbf{X}^{(t)}\mathbf{1}_{n}$, we have $\bar{\mathbf{x}}^{(t+1)}=\bar{\mathbf{x}}^{(t)}-\alpha\bar{\mathbf{m}}^{(t)}$. This implies that although the parameters differ across agents, the EDM algorithm exhibits the same update pattern as standard momentum SGD in the average sense.
@@ 4.1 Auxiliary Variables
To measure the convergence rate, we primarily focus on $f(\bar{\mathbf{x}}^{(t)})$ and its gradient $\nabla f(\bar{\mathbf{x}}^{(t)})$. We begin by defining an auxiliary sequence for $\bar{\mathbf{x}}^{(t)}$, which is commonly used in the analysis of momentum-based SGD methods. Let $\mathbf{z}^{(0)}=\bar{\mathbf{x}}^{(0)}$ and define
@@
The following two lemmas establish the connection between $\mathbf{z}^{(t)}$ and $\mathbf{X}^{(t)}$ under the non-convex cases and PL conditions, respectively.
@@ 4.2 Pseudo Deviation Variance Decomposition
According to Lemma 1 and 2, our next step is to conduct an in-depth investigation of $\mathbb{E}\|\mathbf{X}^{(t)}-\overline{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}$. To simplify the notation, we introduce $\mathbf{P}_{\mathbf{I}}=\mathbf{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$, thus $\|\mathbf{X}^{(t)}-\overline{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}=\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\|_{\mathrm{F}}^{2}$.
The momentum updating structure complicates the exploitation of the independence of $\bm{\Xi}^{(t)}$ in the stochastic gradient. To address this challenge, we introduce the following variables: let $\widetilde{\mathbf{M}}^{(-1)}=\mathbf{0}$ and $\widetilde{\mathbf{M}}^{(t)}=(1-\beta)\sum_{j=0}^{k}\beta^{t-j}\nabla\mathbf{f}(\mathbf{X}^{(j)})$, $\widetilde{\mathbf{X}}^{(-1)}=\widetilde{\mathbf{X}}^{(0)}=\mathbf{X}^{(0)}$ and
@@
By applying Cauchy’s inequality, we obtain:
@@
Suppose that Assumptions 1 and 3 hold. Let $\mathbf{X}^{(t)}$ is the sequence of parameters obtained from Algorithm 1 and $\widetilde{\mathbf{X}}^{(t)}$ be defined as above. Then, we have:
@@
This lemma provides an upper bound for $\mathbb{E}\|\mathbf{P}_{\mathbf{I}}(\mathbf{X}^{(t)}-\widetilde{\mathbf{X}}^{(t)})\|_{\mathrm{F}}^{2}$, which is independent of $\beta$. Thus, we should only focus on the pseudo-deterministic part $\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}$. We call it pseudo-deterministic because it eliminates randomness in the gradients while still retaining randomness in $\mathbf{X}^{(t)}$.
@@ 4.3 Transformations and Consensus Inequalities
Building on the ideas presented in Alghunaim and Yuan (2022), we can observe that $\widetilde{\mathbf{X}}^{(t)}$ exhibits a recursive structure similar to that of the SUDA framework described in their work. Specifically, we define $\widetilde{\mathbf{X}}^{(0)}=\mathbf{X}^{(0)}$, $\widetilde{\mathbf{M}}^{(0)}=(1-\beta)\nabla\mathbf{f}(\mathbf{X}^{(0)})$, $\widetilde{\mathbf{M}}^{(-1)}=\mathbf{0}$, $\widetilde{\mathbf{Y}}^{(0)}=\mathbf{0}$. For all $t\geq 0$, we have:
@@
The following lemma presents a transformation of $\widetilde{\mathbf{X}}^{(t)}$, which is used to derive the upper bound of $\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{X}}^{(t)}\|^{2}$, and consequently for $\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\|^{2}$. This lemma establishes an upper bound on the data consensus deviation that occurs during the iteration, which is central to our analysis of the convergence properties of algorithms.
@@
In this Lemma, we do not specify the exact form of $\widetilde{\mathbf{N}}^{(t)}$. Because taking different values for $\widetilde{\mathbf{N}}^{(t)}$ in different proofs can simplify our proof, even though all forms are intended to measure the difference between the current updates and $\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t)})$. In the subsequent proofs, we will adopt $\widetilde{\mathbf{N}}^{(t)}=(1-\beta)\sum_{j=0}^{t}\beta^{t-j}\nabla\mathbf{f}(\overline{\mathbf{X}}^{(j)})+\beta^{t+1}\nabla\mathbf{f}(\overline{\mathbf{X}}^{(0)})$ for the non-convex case and $\widetilde{\mathbf{N}}^{(t)}=\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t)})$ for the PL condition.
@@ 4.4 Non-Convex
In practical optimization problems, we typically select $\beta\in[0.8,0.99)$, allowing us to treat $\beta$ as a fixed value. When $\beta=0$, the algorithm simplifies to the ED/D2 method. Theorem 5 aligns with the latest conclusions of the ED/D2 framework discussed in Alghunaim and Yuan (2022) with the corresponding bounds illustrated in Table 1. In contrast to their findings, we obtained a bound that is unaffected by the minimum eigenvalue of the communication matrix through a refined scaling of the initial values. Consequently, we achieve a tighter upper bound when $\beta=0$. This results in the momentum-based algorithm exhibiting a convergence rate that is at least comparable to the ED/D2 algorithm. This feature shows the strength of our algorithm.
@@
Unlike Quasi-Global and DecentLaM, which are also the momentum-based variants of distributed stochastic gradient descent, EDM can ultimately eliminate the effects of data heterogeneity. Compared with another kind of bias-correction algorithm (DSGT and DSGT-HB), EDM can eliminate data heterogeneity at a rate of $O(\alpha^{2}(1-\lambda)^{-2}T^{-1})$, while DSGT at a rate of $O(\alpha^{2}(1-\lambda)^{-3}T^{-1})$. This distinction will be demonstrated in our subsequent real-data analysis.
@@ 4.5 PL-Condition
We compare EDM to other momentum-based distributed algorithms under the PL condition. Both the convergence bounds of Quasi-Global and DecentLaM fail to eliminate the effect of heterogeneity. Currently, no work proves the complete convergence properties of the DSGT algorithm with momentum under the Polyak-Lojasiewicz (PL) condition, or even under the strong convexity condition. While Huang et al. (2024) demonstrates convergence properties when $\beta$ is sufficiently large, this result does not extend to the original algorithm (i.e. when $\beta=0$). In contrast, our findings impose no constraints on the choice of $\beta$ and encompass the conclusions of ED/D2. Furthermore, our approach can be generalized to establish convergence properties for momentum-based DSGT algorithms.
On the other hand, Huang et al. (2024) employs the technique of Loopless Chebyshev Acceleration (LCA) to reduce the communication gap in DSGT from $(1-\lambda)^{-1}$ to $(1-\lambda)^{-1/2}$, resulting in a better upper bound on convergence. However, this technique is not exclusive to DSGT; other distributed algorithms can also benefit from LCA to reduce the communication gap. But this aspect is not the focus of our article.